OpenStudy (anonymous):

Simplify: (1/27)^1/3

OpenStudy (anonymous):

1/27^1/3 = 1/(3^3*1/3) = 1/3

OpenStudy (anonymous):

Thank you!!!!!

OpenStudy (anonymous):

Ermm, how would i solve (1/27)^-1/3 ?

OpenStudy (anonymous):

simplify*

OpenStudy (anonymous):

It will be come 3. Negative power is same as positive power of the "inverse".